Write an equation of the cosine function with amplitude 3 and period 4π
1
Expert's answer
2012-02-07T13:04:17-0500
f(x)=3cos(x/2)
Let's verify if written function meets the requirements:
|3cos(x/4)|<=3& -amplitude
f(x+4pi)=2cos((x+4pi)/2)=2cos(x/2+2pi)=2cos(x/2)=f(x) - so 4pi is the period